# BoundMatch: Boundary detection applied to semi-supervised segmentation for urban-driving scenes

### v3

@@ 4.2.3 Pascal VOC 2012
Classic Split: Tab. 3 presents the performance comparison on the Classic splits for both ResNet-50 and ResNet-101 backbones. While SAMTH+BoundMatch improves performance relative to earlier CR methods like AugSeg and UniMatch, it does not reach the state-of-the-art results achieved by the most recent approaches, such as PrevMatch and CW-BASS, on this dataset. We hypothesize this performance gap is partly attributable to the weak CR approach of SAMTH, making it less effective for this dataset. To further assess BoundMatch’s boundary mechanism independently of our SAMTH baseline, we integrated it with UniMatch and PrevMatch again. When combined with these strong baselines, BoundMatch consistently yields mean IoU gains over the original methods, achieving results that are competitive with, or reach, current state-of-the-art performance (Tab. 3).
@@
Blended Split: This split contains potentially even more noisy GT labels, particularly around object boundaries. Tab. 4 shows trends similar to the Classic split. Again, the combinations of BoundMatch with UniMatch or PrevMatch achieve performance competitive with recent SOTA methods. Furthermore, BoundMatch outperforms CFCG [13] in this split, demonstrating its robustness even in the presence of noisy boundary annotations.
@@ 4.2.4 Recent Benchmarks with Transformer Backbones
As shown in Tab. 5, SAMTH+BoundMatch outperforms current state-of-the-art methods, UniMatch-V2 [49] and SegKC [50], across both DINOv2-S and DINOv2-B backbones on the challenging $\nicefrac{{1}}{{16}}$ and $\nicefrac{{1}}{{8}}$ splits. These results demonstrate that BoundMatch effectively scales to modern foundation models while maintaining its boundary-aware advantages. Qualitative visualizations are provided in Appendix E.3.
@@ 4.3 Ablation Studies and Insights
In this section, we conduct ablation studies to analyze the impact of each proposed component and design choice. We report results on the Cityscapes dataset under the $\nicefrac{{1}}{{16}}$ labeled data split (unless explicitly stated), and our evaluation metrics include the mean Intersection-over-Union (IoU), boundary IoU (BIoU), and boundary F1 score (BF1). We also report comparisons with BoundaryMatch, computational costs, and BoundMatch’s effectiveness in real-world semi-supervised learning settings.
@@ 4.3.1 Individual efficacy of the proposed components
Table 6: Component analysis of our framework.
Modules enabled Metrics (%)
BCRM BSF SGF IoU BIoU BF1
$75.05$ $54.75$ $59.28$
$75.91$ $55.25$ $60.22$
$76.21$ $55.95$ $61.97$
$76.10$ $55.70$ $61.64$
76.49 56.09 62.72
Tab. 6 reports the quantitative results for various combinations of our proposed components, starting from the SAMTH baseline which lacks explicit boundary supervision. The addition of our core BCRM framework alone results in notable gains in both segmentation and boundary quality metrics. Further incorporating the BSF module builds upon this, enhancing performance by effectively integrating learned boundary cues into the segmentation head. Finally, adding the SGF module provides complementary benefits; its strategy of refining boundary predictions using spatial gradients derived from the segmentation mask leads to the highest overall performance observed in our ablation. Per-class performance between SAMTH and SAMTH+BoundMatch is provided in Appendix D.
@@ 4.3.2 Analysis of Hyperparameters
Figure 6: Analysis of the hyperparameters introduced by BoundMatch. (a) Varying the boundary threshold $\tau_{bdry}$; (b) Varying the boundary loss weight $\lambda_{bdry}$. For (b), we set $\tau_{bdry}=0.5$. The results are reported on the Cityscapes dataset with $\nicefrac{{1}}{{16}}$ and $\nicefrac{{1}}{{8}}$ split.
@@
Fig. 6 presents the impact of two key hyperparameters introduced by BoundMatch: the boundary threshold $\tau_{bdry}$ and boundary loss weight $\lambda_{bdry}$. For the boundary threshold, we find that $\tau_{bdry}=0.5$ yields optimal performance across both data splits. This moderate threshold balances two factors: boundary predictions typically have low probability values due to the sparsity of boundary pixels, yet setting the threshold too high would exclude valuable boundary information from the consistency regularization process.
@@
Regarding the boundary loss weight $\lambda_{bdry}$, optimal performance is achieved with values around 1.0 for both splits. Interestingly, the $\nicefrac{{1}}{{8}}$ split shows greater robustness to variations in $\lambda_{bdry}$ compared to the $\nicefrac{{1}}{{16}}$ split, suggesting that increased labeled data provides more stable training dynamics that are less sensitive to the precise weighting of the boundary loss.
4.3.3 Binary vs Semantic Boundaries
Table 7: Binary vs. Multi-Label Boundaries.
IoU BIoU BF1
SAMTH $75.05$ $54.75$ $59.28$
+ Binary $75.90$ $55.38$ $60.10$
+ Multi-Label 76.49 56.09 62.72
An important design decision in our method is whether to model boundaries as binary or multi-label (semantic). As shown in Tab. 7, while both formulations improve upon the SAMTH baseline, the semantic boundary formulation achieves the highest improvement in segmentation performance (with IoU increasing from 75.05% to 76.49%). Performance gains from semantic boundaries are supported by experiments in the fully supervised setting for urban scene segmentation [11]. We believe this forces the model to learn not just geometric structures but also semantic context of the boundaries. This suggests that incorporating semantic context into the boundary predictions allows the network to better differentiate between object classes at the boundaries.
@@ 4.3.4 Boundary Evaluation
Figure 7: Qualitative comparison illustrating the impact of Spatial Gradient Fusion (SGF) on boundary predictions (Cityscapes dataset, $\nicefrac{{1}}{{16}}$ split). Boundaries generated using only the BCRM component tend to be thicker and contain more artifacts (false positives). In contrast, the full BoundMatch framework, incorporating the SGF module, produces visibly sharper and cleaner boundaries. SGF achieves this by refining the initial boundary predictions, leveraging spatial gradient information derived from the segmentation mask. This refinement process is crucial for generating higher-quality boundary pseudo-labels, subsequently enhancing the effectiveness of the CR. Best viewed in color and zoomed in.
Fig. 7 presents qualitative boundary predictions at different stages of our framework. The BCRM component alone produces thicker boundaries with more artifacts, while the full BoundMatch framework generates significantly sharper and cleaner boundaries. This visual improvement demonstrates the crucial role of the BSF and SGF modules in refining boundary predictions through their fusion mechanisms.
To quantitatively assess these boundary improvements, we employ two specialized metrics beyond standard mIoU, which is known to be insensitive to boundary-specific errors [10, 63]. We adopt Boundary IoU (BIoU) and Boundary F1 Score (BF1), which explicitly measure boundary alignment and detection quality, respectively (see Appendix C for detailed definitions).
The quantitative results in Tab. 6 and Tab. 7 confirm our qualitative observations. Each boundary-aware component contributes to improved boundary metrics: BCRM alone increases BF1 from 59.28% to 60.22%, while BSF and SGF provide additional gains of approximately 1.7% and 1.4% respectively. Crucially, combining all components yields the highest performance (62.72% BF1), demonstrating their complementary nature. Furthermore, semantic boundaries consistently outperform binary boundaries across all metrics, validating our design choice to use class-specific boundary supervision. These boundary quality improvements directly translate to better segmentation accuracy, particularly for safety-critical applications requiring precise object delineation. Per-class analysis is provided in Appendix D.
@@ 4.3.5 Comparisons with BoundaryMatch
Table 8: Comparison between our BoundMatch and BoundaryMatch [15]. BoundaryMatch is our reproduced results. All models use DeepLabV3$+$ with ResNet-101.
@@
UniMatch + BoundMatch (ours) $76.0$ $77.4$ $58.0$ $64.5$
We compared BoundMatch against BoundaryMatch [15], a recent method that also utilizes boundary CR for SS-SS. For a fair comparison based on our experimental setup, we reimplemented BoundaryMatch and trained it using identical settings. Quantitative results are presented in Tab. 8 which also includes the reported mIoU for BoundaryMatch.
BoundaryMatch achieves higher mIoU compared to UniMatch baseline in Pascal VOC, but does not seem to improve IoU on Cityscapes. UniMatch+BoundMatch achieves +0.9% mIoU and SAMTH+BoundMatch achieves an even higher +1.4% compared to BoundaryMatch on Ciyscapes. Furthermore, boundary metrics (BIoU and BF1) show significant improvements with BoundMatch, indicating that our method is more effective in capturing boundary details, while BoundaryMatch has similar boundary metrics to UniMatch.
We attribute BoundMatch’s superior performance to its more comprehensive handling of boundary information. Specifically, our method employs: (1) semantic boundary supervision, providing richer class-specific edge cues compared to binary boundaries; (2) boundary heads leveraging hierarchical features from the backbone with deep supervision, known to be effective for capturing crucial multi-scale boundary details [26]; and (3) dedicated fusion modules (BSF, SGF) that integrate boundary information back into the segmentation process and refine boundary predictions. In contrast, the BoundaryMatch approach utilizes binary boundaries derived from only a single backbone feature stage. This suggests that simpler boundary consistency regularization might not fully exploit the potential of boundary cues, whereas our multi-faceted approach proves more effective in the SS-SS context.
@@ 4.3.6 Effect of Harmonious BN update strategy
Table 9: Effect of Harmonious Batch-Norm (HBN) update.
@@
SAMTH (+ HBN) $70.7$ +7.9 $75.1$ +5.7
Figure 8: Training curves of AugSeg, CutMix-MT, and SAMTH. The bold lines are method trained with HBN update strategy, while the dotted lines are the ones without.
To further examine the effect of our Harmonious Batch Normalization (HBN) update strategy, we compare different BN update approaches in Tab. 9. Methods like CutMix-MT, ReCo, and U2PL also update the teacher’s BN statistics via forward passes. However, unlike HBN (which processes the complete batch through the teacher specifically for BN updates, as described in Sec. 3.3), these methods often introduce discrepancies between the inputs used for the teacher’s BN statistics and the student’s training inputs. Other approaches, such as iMAS and AugSeg, update the teacher’s BN buffers via EMA based on the student’s statistics. As demonstrated in the table, incorporating our HBN strategy consistently yields significant improvements in segmentation performance across various base methods. For example, applying HBN to SAMTH results in a 5.7% absolute improvement in mean IoU on Cityscapes (from 69.4% to 75.1%). The training curves in Fig. 8 provide further evidence, indicating that not only does HBN results in higher final performance, but it also stabilizes the training process.
@@ 4.3.7 Computational Cost
UniMatch $75.3$ $55.1$ $61.0$ $1.21$s $20.6$GB $191$G $40.5$M $23.4$
SAMTH $75.1$ $54.8$ $59.3$ $0.92$s $9.0$GB $191$G $40.5$M $23.4$
+ BoundMatch $76.5$ $56.1$ $62.7$ $1.59$s $16.5$GB $192$G $40.6$M $19.0$
Tab. 10 summarizes the computational costs of our approach during both training and inference. During training, SAMTH+BoundMatch requires more resources than SAMTH: iteration time increases from 0.92s to 1.59s and memory usage rises from 9.0GB to 16.5GB. The increased iteration time primarily stems from the on-the-fly boundary generation pipeline [11], which computes ground-truth boundaries from labeled data at each iteration. The higher memory usage results from processing spatially large feature maps in the boundary detection head, particularly from early backbone stages. Notably, despite these increases, our total memory footprint remains lower than UniMatch (20.6GB) while iteration time gap reduces to only 0.38s.
For inference, the computational overhead is minimal, with only a modest decrease in FPS (from 23.4 to 19.0) while FLOPs and parameters remain nearly unchanged. These costs are justified by the significant performance gains: +1.4% mIoU overall and up to +3.4% improvement in boundary-specific metrics. The training overhead could be reduced through optimizations of the preprocessing pipeline, which we leave for future work. Given that training is a one-time cost while inference benefits are permanent, this trade-off is particularly favorable for applications requiring precise boundary delineation.

### v4

@@ 4.2.3 Pascal VOC 2012
Classic Split: Tabs.˜4 and 16 presents the performance comparison on the Classic splits for both ResNet-50 and ResNet-101 backbones. While SAMTH+BoundMatch improves performance relative to earlier CR methods like AugSeg and UniMatch, it does not reach the state-of-the-art results achieved by the most recent approaches, such as PrevMatch and CW-BASS, on this dataset. We hypothesize this performance gap is partly attributable to the weak CR approach of SAMTH, making it less effective for this dataset. To further assess BoundMatch’s boundary mechanism independently of our SAMTH baseline, we integrated it with UniMatch and PrevMatch again. When combined with these strong baselines, BoundMatch consistently yields mean IoU gains over the original methods, achieving results that are competitive with, or reach, current state-of-the-art performance (Tab.˜4).
@@
Blended Split: This split contains potentially even more noisy GT labels, particularly around object boundaries. Tabs.˜5 and 17 shows trends similar to the Classic split. Again, the combinations of BoundMatch with UniMatch or PrevMatch achieve performance competitive with recent SOTA methods. Furthermore, BoundMatch outperforms CFCG [13] in this split, demonstrating its robustness even in the presence of noisy boundary annotations.
@@ 4.2.4 Recent Benchmarks with Transformer Backbones
As shown in Tab.˜6, SAMTH+BoundMatch outperforms current state-of-the-art methods, UniMatch-V2 [49] and SegKC [50], across both DINOv2-S and DINOv2-B backbones on the challenging $\nicefrac{{1}}{{16}}$ and $\nicefrac{{1}}{{8}}$ splits. These results demonstrate that BoundMatch scales to modern foundation models while maintaining its boundary-aware advantages. Qualitative visualizations are provided in Appendix E.4.
Table 7: Component analysis of our framework.
ResNet-50 ResNet-101
$\nicefrac{{1}}{{16}}$ $\nicefrac{{1}}{{8}}$ $\nicefrac{{1}}{{16}}$ $\nicefrac{{1}}{{8}}$
BCRM BSF SGF IoU BIoU BF1 IoU BIoU BF1 IoU BIoU BF1 IoU BIoU BF1
$75.1$ $54.8$ $59.3$ $77.3$ $56.8$ $61.8$ $75.5$ $55.8$ $61.4$ $77.9$ $57.9$ $63.1$
$75.9$ $55.3$ $60.2$ $77.6$ $56.9$ $62.3$ $76.9$ $56.9$ $62.1$ $78.2$ $58.5$ $64.5$
$76.2$ $56.0$ $62.0$ $77.5$ $57.2$ $62.8$ $76.1$ $57.3$ $64.5$ $78.2$ $59.0$ $65.0$
$76.1$ $55.7$ $61.6$ $77.6$ $57.0$ $62.4$ $76.6$ $57.0$ $63.9$ $77.9$ $58.8$ $64.9$
76.5 56.1 62.7 78.1 58.1 64.7 77.9 58.3 65.2 79.0 59.4 66.7
@@ 4.3.1 Individual efficacy of the proposed components
To quantify the individual contributions of each proposed component, we conduct systematic ablation experiments evaluating the Boundary Consistency Regularization Module (BCRM), Boundary-Semantic Fusion (BSF), and Spatial Gradient Fusion (SGF) modules. Tab.˜7 reports mIoU, BIoU, and BF1 across multiple evaluation protocols.
The addition of our core BCRM framework alone results in notable gains in both segmentation (0.3–1.4%) and boundary quality metrics (0.1–1.1% for BIoU and 0.5–1.4% for BF1). Further incorporating the BSF module builds upon this, enhancing performance by integrating learned boundary cues into the segmentation head. BCRM with SGF has similar gains, indicating that refining boundary predictions using segmentation features is also beneficial. Finally, using both BSF and SGF modules provide complementary benefits; its strategy of refining boundary predictions using spatial gradients derived from the segmentation mask leads to the highest overall performance across all protocols (0.8–2.4% gains for mIoU).
@@ 4.3.2 Analysis of Hyperparameters
Figure 6: Analysis of the hyperparameters introduced by BoundMatch. (a) Varying the boundary threshold $\tau_{bdry}$; (b) Varying the boundary-loss weight $\lambda_{bdry}$. For (b) we set $\tau_{bdry}=0.5$. Results are on Cityscapes with the $\nicefrac{{1}}{{16}}$ and $\nicefrac{{1}}{{8}}$ splits using the ResNet-50 backbone.
Figure 7: Per-class performance comparison between SAMTH and SAMTH + BoundMatch on the Cityscapes dataset ($\nicefrac{{1}}{{16}}$ split, ResNet-50 backbone). Green values denote improvements, red values denote degradation.
Figure 8: Comparison of semantic boundary predictions. (a) MF (ODS) score of boundary predictions for BCRM versus BoundMatch. (b) Qualitative comparison illustrating the impact of Spatial Gradient Fusion (SGF) on boundary predictions (Cityscapes dataset, $\nicefrac{{1}}{{16}}$ split). Boundaries generated using only the BCRM component tend to be thicker and contain more artifacts (false positives). In contrast, the full BoundMatch framework, incorporating the SGF module, produces visibly sharper and cleaner boundaries. SGF refines the initial boundary predictions by leveraging spatial gradient information derived from the segmentation mask, yielding higher-quality boundary pseudo-labels that enhance the effectiveness of consistency regularization.
@@
Having established the efficacy of individual components, we now examine the sensitivity to key hyperparameters. Fig.˜6 presents the impact of two key hyperparameters introduced by BoundMatch: the boundary threshold $\tau_{bdry}$ and boundary loss weight $\lambda_{bdry}$ on the Cityscapes dataset using ResNet-50 backbone. For the boundary threshold, we find that $\tau_{bdry}=0.5$ yields optimal performance across both data splits. This moderate threshold balances two factors: boundary predictions typically have low probability values due to the sparsity of boundary pixels, yet setting the threshold too high would exclude valuable boundary information from the consistency regularization process.
@@
Regarding the boundary loss weight $\lambda_{bdry}$, optimal performance is achieved with values around $\lambda_{bdry}=1.0$ for both splits. Interestingly, the $\nicefrac{{1}}{{8}}$ split shows greater robustness to variations in $\lambda_{bdry}$ compared to the $\nicefrac{{1}}{{16}}$ split, suggesting that increased labeled data provides more stable training dynamics that are less sensitive to the precise weighting of the boundary loss.
4.3.3 Binary and Semantic Boundaries
Table 8: Binary vs. Multi-Label Boundaries. "Derived" means the model uses boundaries are derived from the predicted pseudo-labels from the segmentation head. "Learned" means predicted boundaries are directly used as the supervision signal.
$\nicefrac{{1}}{{16}}$ $\nicefrac{{1}}{{8}}$
IoU BIoU BF1 IoU BIoU BF1
SAMTH $75.1$ $54.8$ $59.3$ $77.3$ $56.8$ $61.8$
+ Binary (Derived) $75.3$ $54.6$ $59.4$ $77.3$ $56.9$ $62.0$
+ Binary (Learned) $75.9$ $55.4$ $60.1$ $77.6$ $57.3$ $62.5$
+ Multi-Label (Derived) $75.8$ $55.4$ $60.8$ $77.4$ $57.1$ $62.5$
+ Multi-Label (Learned) 76.5 56.1 62.7 78.1 58.1 64.7
An important design decision in our method is what types of boundaries should be used: derived versus learned, and binary versus semantic (multi-label). As shown in Tab.˜8, using learned multi-label boundaries obtains the highest improvements across both evaluation protocols, outperforming baseline with binary boundaries and derived boundaries from segmentation pseudo-labels. For $\nicefrac{{1}}{{16}}$, IoU improves by $1.4\%$, BIoU by $1.3\%$, and BF1 by $3.4\%$, indicating that learned class-aware boundary supervision provides benefits over the baseline approach.
Using derived boundaries show minimal improvements over SAMTH compared to using learned boundaries for consistency regularization. We believe this is because generating boundaries (e.g. via Laplacian operators) from often noisy segmentation pseudo-labels propagates these noise to the boundary predictions, limiting their effectiveness. In contrast, learning boundaries directly from dedicated boundary predictions allows the model to focus on the geometric structures of the boundaries independent of the segmentation task, and the ability to threshold boundaries by confidence ($\tau_{bdry}$) allows the model to learn more robust boundary representations. Furthermore, our bidirectional fusion allows the boundary and segmentation heads to mutually benefit from each other, where SGF module enables learning with segmentation mask cues.
4.3.4 Per-Class Performance
To understand how these boundary formulations affect different object categories, we analyze per-class performance. Here we present the per-class performance comparison between SAMTH and BoundMatch for Cityscapes $\nicefrac{{1}}{{16}}$ split in LABEL:fig:per_class_iou, LABEL:fig:per_class_biou and LABEL:fig:per_class_bf1 for IoU, BIoU, and BF1 metrics respectively. For IoU, we observe that BoundMatch improves the performance on most of the classes, especially on vehicles (e.g. car, truck, bus, and train) and people (i.e. person and rider). While thin and small objects (e.g. traffic light and traffic sign) also see improvements, SAMTH+BoundMatch seems to have less impact on pole although we see qualitative improvements for poles in LABEL:fig:app_successful_cases This may be due to pixel imbalance issues, as poles occupy a very small fraction of the image area, but it could also be caused by noisy annotations in the ground-truth which amplifies false positives as shown in LABEL:fig:app_failure_cases due to BoundMatch focusing more on the details. The fence category presents an interesting exception with decreased IoU but maintained boundary metrics, likely due to annotation ambiguities where fences are labeled as solid regions despite their sparse structure as shown in Fig.˜10. BoundMatch improves BIoU and BF1 metrics for most object categories, though minor degradations are observed for certain classes with ambiguous boundaries (e.g. fence, terrain), which aligns with the challenges discussed in our limitations.
@@ 4.3.5 Boundary Evaluation
Fig.˜8 presents quantitative and qualitative results comparing boundary predictions from BCRM alone versus the full BoundMatch framework (BCRM+BSF+SGF). Looking at the mean F-measure (MF) at optimal dataset scale (ODS), we observe that BoundMatch consistently outperforms BCRM alone across various experiment protocols, indicating that the SGF refinement module improves boundary quality. Qualitatively, BCRM alone produces thicker boundaries with more artifacts, while the full BoundMatch framework generates sharper and cleaner boundaries. This improvement demonstrates the benefit of SGF’s spatial gradient fusion in refining boundary predictions for more reliable pseudo-labels.
@@ 4.3.6 Comparisons with BoundaryMatch
Table 9: Comparison between our BoundMatch and BoundaryMatch [15]. BoundaryMatch is our reproduced results. All models use DeepLabV3++ with ResNet-101.
@@
We compared BoundMatch against BoundaryMatch [15], a recent method that also utilizes boundary CR for SS-SS. For a fair comparison based on our experimental setup, we reimplemented BoundaryMatch and trained it using identical settings. Quantitative results on Pascal VOC and Cityscapes dataset using ResNet-101 backbone are presented in Tab.˜9 which also includes the reported mIoU for BoundaryMatch.
BoundaryMatch achieves higher mIoU compared to UniMatch baseline in Pascal VOC, but does not seem to improve IoU on Cityscapes. With multi-label boundaries, the metrics does surpass UniMatch slightly. UniMatch+BoundMatch achieves 0.9% mIoU and SAMTH+BoundMatch achieves an even higher 1.4% compared to BoundaryMatch on Cityscapes. Furthermore, boundary metrics (BIoU and BF1) also show notable improvements with BoundMatch, indicating that our method is more effective in capturing boundary details, while BoundaryMatch has similar boundary metrics to UniMatch.
BoundMatch’s improved performance stems from four key technical advantages that address the limitations of simpler consistency approaches: (1) semantic boundary supervision, providing richer class-specific edge cues compared to binary boundaries; (2) the use of learned boundaries as pseudo-labels for consistency regularization; (3) boundary heads leveraging hierarchical features from the backbone with deep supervision, known to be effective for capturing multi-scale boundary details [26]; and (4) dedicated fusion modules (BSF, SGF) that integrate boundary information back into the segmentation process and refine boundary predictions. This suggests that simpler boundary consistency regularization might not fully exploit the potential of boundary cues, whereas our multi-faceted approach proves more effective in the SS-SS context.
@@ 4.3.7 Effect of Harmonious BN update strategy
Table 10: Effect of Harmonious Batch-Norm (HBN) update on Pascal VOC (92 labeled image) and Cityscapes ($\nicefrac{{1}}{{16}}$ split) datasets using ResNet-50 backbone. We report the mean IoU and the absolute gains ($\Delta$) after applying HBN to various base methods.
@@
SAMTH $70.7$ +7.9 $75.1$ +5.7
Figure 9: Training curves of AugSeg, CutMix-MT, and SAMTH on the Cityscapes $\nicefrac{{1}}{{16}}$ benchmark using ResNet-50 backbone. The bold lines are method trained with HBN update strategy, while the dotted lines are the ones without.
To further examine the effect of our Harmonious Batch Normalization (HBN) update strategy, we compare different BN update approaches in Tab.˜10. Methods like CutMix-MT, ReCo, and U2PL also update the teacher’s BN statistics via forward passes. However, unlike HBN (which processes the complete batch through the teacher specifically for BN updates, as described in Sec.˜3.3), these methods often introduce discrepancies between the inputs used for the teacher’s BN statistics and the student’s training inputs. Other approaches, such as iMAS and AugSeg, update the teacher’s BN buffers via EMA based on the student’s statistics. In our experiments, incorporating HBN improves segmentation performance for most tested baseline methods, with mIoU gains ranging from 0.8–16.4%. For example, applying HBN to SAMTH results in a 5.7% absolute improvement in mean IoU on Cityscapes (from 69.4% to 75.1%). The training curves in Fig.˜9 provide further evidence, indicating that not only does HBN results in higher final performance, but it also stabilizes the training process.
@@ 4.3.8 Computational Cost
UniMatch $75.3$ $55.1$ $61.0$ $1.21$s 20.620.6GB 191191G 40.540.5M $23.4$
PrevMatch $75.7$ $55.3$ $60.7$ $1.56$s 21.421.4GB 191191G 40.540.5M $23.4$
SAMTH $75.1$ $54.8$ $59.3$ $0.92$s 9.09.0GB 191191G 40.540.5M $23.4$
+ BCRM $75.9$ $55.3$ $60.2$ $1.24$s 13.213.2GB 191191G 40.540.5M $23.4$
+ BCRM + BSF $76.2$ $56.0$ $62.0$ $1.32$s 13.513.5GB 192192G 40.640.6M $19.0$
+ BCRM + SGF $76.1$ $55.7$ $61.6$ $1.45$s 15.015.0GB 191191G 40.540.5M $23.4$
+ BoundMatch $76.5$ $56.1$ $62.7$ $1.59$s 16.516.5GB 192192G 40.640.6M $19.0$
Tab.˜11 analyzes the computational overhead of BoundMatch’s individual components. BCRM adds 0.32s and 4.2GB for multi-scale boundary processing, BSF contributes minimal overhead (0.08s, 0.3GB), while SGF requires 0.21s and 1.8GB for gradient computations. The complete framework efficiently combines these modules with shared computations, resulting in a total training overhead of 0.67s and 7.5GB compared to SAMTH baseline, primarily due to the on-the-fly boundary generation pipeline [11] and the fusion modules. Despite this increase, BoundMatch remains more memory-efficient than UniMatch (20.6GB) and have similar per-iteration time as PrevMatch (1.56s) while achieving accuracy gains. Optionally, if training costs are a concern, one can disable BSF and only use BCRM+SGF which still provides 1.0% mIoU improvement over SAMTH with only 0.53s and 6.0GB overhead.
At inference, BoundMatch introduces minimal overhead (1G FLOPs, 0.1M parameters), reducing FPS from 23.4 to 19.0 for 1.4% mIoU and up to 3.4% boundary metric improvements. BSF slightly increases FLOPs and reduces FPS due to additional convolutions required feature fusion, thus if the trade-off is not desired, one can use the BCRM+SGF combination which has no impact on inference cost. BoundMatch’s modular design allows users to change feature fusion modules entirely (e.g. use attention-based fusion), which may offer different trade-offs between accuracy and efficiency. Investigations into better feature fusion strategies that are computationally efficient are left for future work.